# Supplementary Materials

### v1

@@ Appendix B Additional Results and Complete Proofs
Lemma B.1.
@@
where $\beta$ is a scaling factor ensuring that the Euclidean norms of $Sx$ and $\beta e_{1}$ are equal.
@@
Next, we show that $A\in G$. Consider $A^{T}DA$:
$A^{T}DA$ $=\mathopen{}\left({S^{-1}RS}\,\negthickspace\right)\mathclose{}^{T}D\mathopen{}\left({S^{-1}RS}\,\negthickspace\right)\mathclose{}$
$=S^{T}R^{T}\mathopen{}\left({S^{-1}DS^{-1}}\,\negthickspace\right)\mathclose{}RS$
$=S^{T}R^{T}IRS$
$=S^{T}IS$
$=D.$ (21)
Thus, $A\in G$, as required. ∎
Remark B.1.
The Lemma. B.1 is also applicable when $d_{i}<0,\;\forall i\in[n]$. We just need to appropriately modify the matrix $S$.
Lemma B.2.
@@
where $d_{1},\dots,d_{k-1}>0$, and $E_{i,j}=e_{i}e_{j}^{T}$ with $e_{i}$ denoting the standard basis vectors in $\mathbb{R}^{n}$. Consider a vector $y\in\mathbb{R}^{n}$ of the form:
@@
Step 2: Verify $Ay=ae_{1}$.
Substituting $y=ae_{1}+be_{k}$ into $Ay$, we have:
$Ay$ $=\mathopen{}\left({E_{1,1}+d_{1}abE_{k,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({ae_{1}+be_{k}}\,\negthickspace\right)\mathclose{}$
$=ae_{1}+d_{1}a^{2}be_{k}-d_{1}a^{2}be_{k}$
$=ae_{1}.$ (30)
Thus, $Ay=ae_{1}$, as required. Conclusion: We have constructed $A\in G$ such that $Ay=ae_{1}$. Therefore, the lemma is proved. ∎
Remark B.2.
The Lemma. B.2 is also applicable when $d_{i}<0,\;\forall i\in\{1,\dots,k\}$. We just need to appropriately modify the matrices $S$ and $R$.
Lemma B.3.
@@
Using the Lemma. B.1, B.2 and B.3, we provide the complete proof of Theorem. 4.1. See 4.1
@@
Apply Lemma B.1 to $\mathopen{}\left({D_{p},x_{p}}\,\negthickspace\right)\mathclose{}$ and $\mathopen{}\left({D_{n},x_{n}}\,\negthickspace\right)\mathclose{}$, yielding matrices $A_{p}$ and $A_{n}$ such that:
@@
Applying Lemma B.2 to $\mathopen{}\left({D_{nz},x_{nz}}\,\negthickspace\right)\mathclose{}$, we find $A_{nz}$ such that:
@@
Applying Lemma B.3 to $\mathopen{}\left({D,A_{2}A_{1}x}\,\negthickspace\right)\mathclose{}$ (for the case of $x^{T}Dx>0$), we find $A_{3}$ such that:
@@ Conclusion
From Theorem 4.2, we know that $\phi_{s}(.)$ belongs to the group $G$. Thus, we can take the inverse (modulo the stabilizer subgroup of $x$) of $\phi_{s}$ and express the equation as follows:
$\mathopen{}\left({\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}}\,\negthickspace\right)\mathclose{}^{-1}f(x)=\phi_{n}\mathopen{}\left({\|x\|_{A}}\,\negthickspace\right)\mathclose{}$ (54)
If $A$ is either a positive definite or a negative definite matrix, the map $x\mapsto\frac{x}{\|x\|_{A}}$ is smooth on the domain $\mathbb{R}^{n}\setminus U$. Since the group action is smooth, we conclude that the map $x\mapsto\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}^{-1}$ is smooth. Therefore, if $f$ is smooth, $\phi_{n}$ must also be smooth.
In the case where $A$ is neither positive definite nor negative definite, the map $x\mapsto\frac{x}{\|x\|_{A}}$ is not smooth, but it is continuous. Hence, if $f$ is continuous, $\phi_{n}$ is continuous as well. ∎
Theorem B.4 (Classification of $p$-tuples up to the $A$-orthogonal group, cf. [Jacobson, 1985, Chapter 6]).
] Let $A$ be a non-degenerate real symmetric $n\times n$ matrix, and consider the bilinear form $\langle x,y\rangle_{A}=x^{\top}A\,y$ on $\mathbb{R}^{n}$. Suppose $(x_{1},\dots,x_{p})$ and $(y_{1},\dots,y_{p})$ are two ordered $p$-tuples in $\mathbb{R}^{n}$. Then the following are equivalent:
@@
In the proof of Theorem 4.2, we utilized the orbit correspondence result (Theorem 4.1) to establish the norm invariance part ($\phi_{n}$) of the function decomposition. In the setting of the diagonal action on tuples, the invariance part ($\phi_{n}$) of the decomposition follows immediately from Theorem B.4. The remainder of the proof proceeds analogously to that of Theorem 4.2. ∎

### v2

@@ Appendix C Additional Results and Complete Proofs
C.1 Canonical Alignment in the Positive Definite Case
We begin with the positive definite case, where the quadratic form $x^{\top}Dx$ defines a weighted Euclidean geometry determined by the positive diagonal entries of $D$. In this setting, all directions are comparable through the group of transformations that preserve this weighted norm. The goal of this subsection is to establish a canonical alignment result showing that any vector can be mapped to a fixed reference direction through an appropriate transformation in the orthogonal group $G=\{A\in\mathbb{R}^{n\times n}:A^{\top}DA=D\}$. This alignment construction will serve as a fundamental building block in the proof of Theorem 4.1, which characterizes the orbits of this group and their correspondence with quadratic-form norms.
Lemma C.1.
@@ C.1 Canonical Alignment in the Positive Definite Case
where $\beta$ is a scaling factor that ensures that the Euclidean norms of $Sx$ and $\beta e_{1}$ are equal.
@@
Next, we show that $A\in G$. Consider $A^{T}DA$. Since $S$ is a diagonal matrix, $S^{T}=S$ and $(S^{-1})^{T}=S^{-1}$.
$A^{T}DA$ $=(S^{-1}RS)^{T}D(S^{-1}RS)$
$=(S^{T}R^{T}(S^{-1})^{T})D(S^{-1}RS)$
$=(SR^{T}S^{-1})D(S^{-1}RS)$ $\text{Since }S^{T}=S$
$=SR^{T}(S^{-1}DS^{-1})RS$ By associativity
$=SR^{T}IRS$ $\text{Since }S=\sqrt{D},S^{-1}DS^{-1}=I$
$=SR^{T}RS$ $\text{Since }R\in O(n),R^{T}R=I$
$=SIS=S^{2}=D.$
This shows conclusively that $A\in G$. ∎
Remark C.1.
The Lemma. C.1 is also applicable when $d_{i}<0,\;\forall i\in[n]$. We just need to appropriately modify the matrix $S$.
Lemma C.1 thus guarantees that, in the positive definite setting, every orbit under the $D$-orthogonal group admits a canonical representative obtained by aligning an arbitrary vector with the first coordinate axis. This serves as the foundational case for proving Theorem 4.1; the subsequent subsections extend this reasoning to semi-definite and indefinite quadratic forms, where certain directions may lie on null subspaces or carry opposite signs.
C.2 Canonical Alignment in the Positive Semi-Definite Case
Lemma C.2.
@@ C.2 Canonical Alignment in the Positive Semi-Definite Case
where $d_{1},\dots,d_{k-1}>0$, with $k\neq 1$, and $E_{i,j}=e_{i}e_{j}^{T}$ with $e_{i}$ denoting the standard basis vectors in $\mathbb{R}^{n}$. Consider a vector $y\in\mathbb{R}^{n}$ of the form:
@@
Step 2: Verify $Ay=ae_{1}$
We are given the vector $y=ae_{1}+be_{k}$. Since $k\neq 1$, ensuring that $e_{1}$ and $e_{k}$ are distinct standard basis vectors. Our goal is to verify that for the constructed matrix $A$, the transformation yields $Ay=ae_{1}$.
The matrix $A$ is given by:
$A=E_{1,1}+d_{1}abE_{k,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],i\neq 1,i\neq k}E_{i,i}$
Let’s apply $A$ to $y$ by distributing the terms:
$Ay=\left(E_{1,1}+d_{1}abE_{k,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],i\neq 1,i\neq k}E_{i,i}\right)(ae_{1}+be_{k})$
We evaluate the effect of each component of $A$ on $y$, noting that the elementary matrix $E_{i,j}$ acts on a basis vector $e_{m}$ as $E_{i,j}e_{m}=\delta_{jm}e_{i}$, where $\delta_{jm}$ is the Kronecker delta.
  1. 1.
    First Term: $E_{1,1}(ae_{1}+be_{k})=aE_{1,1}e_{1}+bE_{1,1}e_{k}=ae_{1}+0=\boldsymbol{ae_{1}}$.
  2. 2.
    Second Term: $d_{1}abE_{k,1}(ae_{1}+be_{k})=ad_{1}abE_{k,1}e_{1}+bd_{1}abE_{k,1}e_{k}=ad_{1}abe_{k}+0=\boldsymbol{d_{1}a^{2}be_{k}}$.
  3. 3.
    Third Term: $d_{1}a^{2}E_{k,k}(ae_{1}+be_{k})=-ad_{1}a^{2}E_{k,k}e_{1}-bd_{1}a^{2}E_{k,k}e_{k}=0-bd_{1}a^{2}e_{k}=\boldsymbol{-d_{1}a^{2}be_{k}}$.
  4. 4.
    Fourth Term: The sum $\sum_{i\in[n],i\neq 1,i\neq k}E_{i,i}$ acts as the identity on basis vectors with indices other than $1$ and $k$. Therefore, its product with $(ae_{1}+be_{k})$ is zero.
Summing the results from each component:
$Ay$ $=\boldsymbol{ae_{1}}+\boldsymbol{d_{1}a^{2}be_{k}}-\boldsymbol{d_{1}a^{2}be_{k}}$
$=ae_{1}.$
Thus, we have verified that $Ay=ae_{1}$, as required.
Conclusion: We have constructed $A\in G$ such that $Ay=ae_{1}$. Therefore, the lemma is proved.
Remark C.2.
The Lemma. C.2 is also applicable when $d_{i}<0,\;\forall i\in\{1,\dots,k\}$. We just need to appropriately modify the matrices $S$ and $R$.
C.3 Canonical Alignment in the Indefinite Case
Lemma C.3.
@@ C.3 Canonical Alignment in the Indefinite Case
C.4 Orbit Space and Quadratic-Form Norm Correspondence
Building upon Lemmas C.1, C.2, and C.3, we now establish Theorem 4.1, which characterizes the structure of group orbits under quadratic-form–preserving transformations. See 4.1
@@ C.4 Orbit Space and Quadratic-Form Norm Correspondence
Apply Lemma C.1 to $\mathopen{}\left({D_{p},x_{p}}\,\negthickspace\right)\mathclose{}$ and $\mathopen{}\left({D_{n},x_{n}}\,\negthickspace\right)\mathclose{}$, yielding matrices $A_{p}$ and $A_{n}$ such that:
@@
Applying Lemma C.2 to $\mathopen{}\left({D_{nz},x_{nz}}\,\negthickspace\right)\mathclose{}$, we find $A_{nz}$ such that:
@@
Applying Lemma C.3 to $\mathopen{}\left({D,A_{2}A_{1}x}\,\negthickspace\right)\mathclose{}$ (for the case of $x^{T}Dx>0$), we find $A_{3}$ such that:
@@ Conclusion
C.5 Decomposition into Scale- and Norm-Invariant Components
Building upon the orbit characterization established in Theorem 4.1, we now derive our main result: a canonical decomposition of any $G$-equivariant function into scale-invariant and norm-invariant components. The orbit-space analysis reveals that the orbit of each input vector is uniquely determined by its quadratic-form norm. In the following result, we further show that the combination of this norm and the corresponding normalized direction completely characterizes each input, thereby enabling the desired factorization of equivariant functions.
The following theorem formalizes this decomposition result.
@@ C.6 Regularity conditions
The proof relies on analyzing the regularity of the components in the rearranged decomposition from Theorem 4.2.
  1. 1.
    From the proof of Theorem 4.2, we can isolate the norm-invariant component $\phi_{n}$ by applying the inverse of the group element returned by $\phi_{s}$.
    $\phi_{n}(\|x\|_{A})=\left(\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\right)^{-1}f(x)$ (57)
    This equation shows that the regularity of $\phi_{n}$ depends on the regularity of $f$ and the map $x\mapsto\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)$. Since $\phi_{s}$ maps to the Lie group $G$, and both group action and group inversion are smooth operations, the regularity of the expression is determined by the map $x\mapsto\frac{x}{\|x\|_{A}}$ and the function $f(x)$ itself.
  2. 2.
    Case 1: A is Positive or Negative Definite
    If matrix $A$ is positive or negative definite, the quadratic form $x^{T}Ax$ is non-zero for all non-zero $x\in\mathbb{R}^{n}$. This means the set $U:=\{x\in\mathbb{R}^{n}:x^{T}Ax=0\}$ contains only the zero vector. The domain of analysis is $\mathbb{R}^{n}\setminus U=\mathbb{R}^{n}\setminus\{0\}$.
    The norm is defined as $\|x\|_{A}=\text{sign}(x^{T}Ax)\sqrt{|x^{T}Ax|}$. When $A$ is definite, $\text{sign}(x^{T}Ax)$ is constant for all $x\neq 0$. The expression inside the square root, $|x^{T}Ax|$, is a smooth and strictly positive polynomial for $x\neq 0$. The square root of a smooth, strictly positive function is smooth.
    Therefore, the map $x\mapsto\|x\|_{A}$ is smooth on $\mathbb{R}^{n}\setminus\{0\}$. Consequently, the map $x\mapsto\frac{x}{\|x\|_{A}}$ is a composition of smooth functions and is itself smooth[cite: 622]. Since the group action is smooth, the function $x\mapsto\left(\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\right)^{-1}$ is also smooth.
    If we assume $f$ is smooth ($C^{\infty}$), then the right-hand side of Eq. (57) is a product of smooth functions, which is smooth. Thus, $\phi_{n}$ must also be smooth[cite: 624].
  3. 3.
    Case 2: A is Indefinite
    If $A$ is neither positive nor negative definite, the set $U$ contains non-zero vectors. The function $g(z)=\sqrt{|z|}$ is continuous everywhere but not differentiable at $z=0$. Because $x^{T}Ax$ can be zero for non-zero $x$, the map $x\mapsto\|x\|_{A}$ is continuous on $\mathbb{R}^{n}$ but is not differentiable on the set $U$. The map $x\mapsto\frac{x}{\|x\|_{A}}$ is therefore continuous on its domain $\mathbb{R}^{n}\setminus U$.
    If we assume $f$ is continuous ($C^{0}$), then the right-hand side of Eq. (57), being a product and composition of continuous functions, is continuous. It follows that $\phi_{n}$ must also be continuous.
This covers both cases and completes the proof. ∎
C.7 Extension to Diagonal action
Theorem C.4 (Classification of $p$-tuples up to the $A$-orthogonal group, cf. (Jacobson, 1985, Chapter 6)).
Let $A$ be a non-degenerate real symmetric $n\times n$ matrix, and consider the bilinear form $\langle x,y\rangle_{A}=x^{\top}A\,y$ on $\mathbb{R}^{n}$. Suppose $(x_{1},\dots,x_{p})$ and $(y_{1},\dots,y_{p})$ are two ordered $p$-tuples in $\mathbb{R}^{n}$. Then the following are equivalent:
@@ C.7 Extension to Diagonal action
In the proof of Theorem 4.2, we utilized the orbit correspondence result (Theorem 4.1) to establish the norm invariance part ($\phi_{n}$) of the function decomposition. In the setting of the diagonal action on tuples, the invariance part ($\phi_{n}$) of the decomposition follows immediately from Theorem C.4. The remainder of the proof proceeds analogously to that of Theorem 4.2. ∎
C.8 Householder reflection
See 5.1
Proof.
The proof combines the canonical decomposition from Theorem 4.2 with the specific properties of the group action and the scale-invariant component.
  1. 1.
    Canonical Decomposition: According to Theorem 4.2, any G-equivariant function $f:X\to Y$ can be written as the action of a scale-invariant component on a norm-invariant component:
    $f(x)=\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\cdot\phi_{n}(\|x\|_{A})$ (58)
    where $\phi_{s}:X\to G$ and $\phi_{n}:\mathbb{R}\to Y$. The symbol $\cdot$‘ denotes the group action of $G$ on the output space $Y$.
  2. 2.
    Group Action as Conjugation: The proposition states that the group action on the output is conjugation. For a group element $g\in G$ and an output object $Y\in Y$, this action is defined as $g\cdot Y=gYg^{-1}$. Applying this definition to our decomposition in Eq. (58), the action of the group element $\phi_{s}(\dots)$ on the output $\phi_{n}(\dots)$ is:
    $f(x)=\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\phi_{n}(\|x\|_{A})\left(\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\right)^{-1}$ (59)
  3. 3.
    Householder Reflection Property: As discussed in Section 5.2, the scale-invariant component $\phi_{s}(\frac{x}{\|x\|_{A}})$ is equivalent to a Householder reflection. A key property of a Householder reflection matrix $R$ is that it is its own inverse, i.e., $R=R^{-1}$. Therefore, we have:
    $\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)=\left(\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\right)^{-1}$ (60)
  4. 4.
    Final Form: By substituting the property from Eq. (60) into Eq. (59), we replace the inverse term with the term itself:
    $f(x)$ $=\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\phi_{n}(\|x\|_{A})\left(\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\right)^{-1}$
    $=\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)\phi_{n}(\|x\|_{A})\phi_{s}\left(\frac{x}{\|x\|_{A}}\right)$
This yields the symmetric form of the function as stated in the proposition. ∎